OpenStudy (anonymous):

how do i graph y-2x≥1

OpenStudy (anonymous):

graph the line y-2x=1 and then shade one side area of that line

OpenStudy (anonymous):

i know, but i forgot how to graph itt..

OpenStudy (anonymous):

Put x=0 and find y1 that will be one point (0,y1) then put y=0 and find x2 that will be second point (x2,0). put these points on the graph and join them

OpenStudy (anonymous):

uhhhhhhhhh):???

OpenStudy (anonymous):

in the equation y-2x=1, replace x with 0 and find y?

OpenStudy (anonymous):

y=2x+1??

OpenStudy (anonymous):

replace x with 0.

OpenStudy (anonymous):

merr.

OpenStudy (anonymous):

y = 2x0 + 1=?

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

That's right. So with x as 0 one of your point is (0,1). Can you plot this point on graph.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Do that, you have one point. Now replace y with 0 and find a value for x.